# From molecules to scaffolds to functional groups: building context-dependent molecular representation via multi-channel learning

### v2

@@ Abstract
Reliable molecular property prediction is essential for various scientific endeavors and industrial applications, such as drug discovery. However, the data scarcity, combined with the highly non-linear causal relationships between physicochemical and biological properties and conventional molecular featurization schemes, complicates the development of robust molecular machine learning models. Self-supervised learning (SSL) has emerged as a popular solution, utilizing large-scale, unannotated molecular data to learn a foundational representation of chemical space that might be advantageous for downstream tasks. Yet, existing molecular SSL methods largely overlook chemical knowledge, including molecular structure similarity, scaffold composition, and the context-dependent aspects of molecular properties when operating over the chemical space. They also struggle to learn the subtle variations in structure-activity relationship. This paper introduces a novel pre-training framework that learns robust and generalizable chemical knowledge. It leverages the structural hierarchy within the molecule, embeds them through distinct pre-training tasks across channels, and aggregates channel information in a task-specific manner during fine-tuning. Our approach demonstrates competitive performance across various molecular property benchmarks and offers strong advantages in particularly challenging yet ubiquitous scenarios like activity cliffs.
1 Introduction
@@
To approach the aforementioned challenges, we introduce a prompt-guided multi-channel learning framework for molecular representation learning. Each of the channels, guided by a specific prompt token, is responsible for learning one dedicated SSL task. Essentially, the pre-trained model is able to learn distinct representation spaces. During fine-tuning, a prompt selection module aggregates representations into a composite representation and uses it for the downstream molecular property predictions. This involves determining which information channel is most relevant to the current application, thereby making the representation context-dependent. We later show how this composite formulation is more resilient to label overfitting and manifests better robustness. In addition, we design the pre-train tasks to form an interpolation from a global view to a local view of the molecular structures. Besides leveraging the global molecule contrastive learning and the local context prediction [38], we introduce a novel task called the scaffold contrastive distancing, highlighting the fundamental role of scaffolds in affecting molecular characteristics and behaviors. Since scaffolds are often treated as starting points for new compound design, scaffold distancing aims to map molecules with similar scaffolds (generated via scaffold-invariant perturbations) closer in the representation space. Additionally, it pushes molecules with different scaffolds apart, where the distance margin is computed adaptively based on structure composition difference. Note that scaffold distancing tackles the partial but core view of the molecules. The overall framework is pre-trained using ZINC15 [42], and evaluated on 7 molecular property prediction tasks in MoleculeNet [56] and 30 binding potency prediction tasks in MoleculeACE [50]. Learning to leverage information from different channels for different applications, our method surpasses various representation learning baselines in both benchmarks. More importantly, our method is shown to handle the challenge of activity cliffs more effectively, whereas competing approaches are more susceptible to the negative transfer, leading to a substantial performance decline. This suggests that these methods may rely more on surface-level patterns even after pre-training or are more susceptible to knowledge forgetting during fine-tuning, causing them to struggle with challenging problems that require a nuanced understanding of chemical knowledge.
Besides reporting the performance metrics, we further examine the shift in representation space during fine-tuning via epoch probing, and discover that our learned representation is better at preserving pre-trained knowledge with better transferability and robustness compared to other baselines. Case study shows that our method has the potential to identify crucial patterns that contribute to activity cliffs even using merely the topological information, such as the molecular graph. We also demonstrate that the learned channel-wise representation distances are highly correlated with the conventional fingerprint/descriptor similarity. These representations can form an implicit representation hierarchy, visualized by iteratively refining the chemical space clustering.
2 Results
2.1 Framework overview
@@
The proposed framework (Figure 1) comprises three novel components compared to the conventional pretrain-finetune paradigm on molecules: (1) The prompt-guided multi-channel learning, (2) contrastive learning with adaptive margin, and (3) scaffold-invariant molecule perturbation.
@@
We now briefly introduce the three learning channels. (i) Molecule distancing (Figure 1b) is archived using a variant of triplet contrastive loss [40]. It considers triplet of molecule samples {anchor, positive, negative}, where negative (i.e., dissimilar) samples are pushed apart against the anchor and positive (i.e., similar) samples by a margin . On top of this, we propose the adaptive margin , as detailed below. It introduces another level of distancing constraints based on the structural similarity of molecule composition. We follow the work of Molecular Contrastive Learning of Representations (MolCLR) [52] and generate positive samples via molecule subgraph masking. (ii) Scaffold distancing (Figure 1c) is a novel contrastive learning task that focuses on scaffold differences. Molecule scaffolds are viewed as the foundation for a range of biologically active molecules. They play a crucial role in drug discovery and medicinal chemistry by providing a starting point for compound designs with desired pharmacological properties. In other words, molecules with similar scaffolds are more likely to possess similar physical (e.g., solubility, lipophilicity) and biological (e.g., conformational property when interacting with a protein) characteristics, and thereby sharing similar semantics. Scaffold distancing contrasts the scaffold-invariant molecule perturbations, as detailed below, against molecules with different scaffolds using the adaptive margin loss. (iii) Context prediction (Figure 1d) involves masked subgraph prediction and motif prediction, which are also adopted in GROVER [38]. For each molecular graph, a random subgraph (i.e., a center atom and its one-hop neighbors) is masked out, and the model aims to reconstruct the subgraph based on its surrounding structures. Motif prediction aims to predict the existence of functional groups within the molecule. This learning channel mainly focuses on the local view of the molecule by identifying the existence of substructure and functional groups, while molecule distancing and scaffold distancing focus more on the global view and partial view, respectively. Each channel has its own readout layer, termed prompt-guided node aggregation (Figure 1e). Presumably, the distribution of atom importance should be different across channels. To further improve the robustness of the learned knowledge under this framework, we incorporate two regularization tricks on the intra-channel node aggregation and the inter-channel alignment. For the molecule and scaffold channels, the aggregation attentions are more encouraged to span across all atoms and scaffold atoms, respectively. Meanwhile, a small set of supervised tasks is utilized to regularize the composite representation under three prompt weight presets, hence improving the channel alignment.
@@
In the fine-tuning stage, the model is initialized with the same molecule encoder and prompt-guided aggregation modules, along with the learned parameters from pre-training. The parameters in the aggregation modules are fixed during fine-tuning as we aim to use it as a pooling layer independent of the downstream applications. In addition, a prompt-tuning module is introduced to determine which channel is most relevant to the current application. It essentially learns a task-specific prompt distribution. The prompt weights are utilized to linearly combine the channel-wise information into a composite molecule representation, which is then used for the task-specific prediction. We initialize the prompt weights by choosing the candidate which leads to the smoothest quantitative structure-property landscape (i.e., lowest roughness index) [2] of the composite representation. More details are included in the Method section.
@@
We further introduce the adaptive margin loss, a variant of the triplet loss [40], that supports the contrastive learning in the first two channels (molecule distancing and scaffold distancing). In the conventional triplet loss, the representation distance between the anchor and negative (i.e., semantic-dissimilar) sample needs to be at least by margin larger than the distance between the anchor and positive (i.e., semantic similar) sample . Note that this margin remains the same for any triplet considered. However, when applied to molecule triplets, it neglects the known structural relationship between molecules (e.g., co-existence of functional groups). To learn a more fine-grained molecule representation space, we propose to adaptively compute the molecule triplet margin based on the Tanimoto similarity between molecule fingerprints. As shown in Figure 1b and c, the adaptive margin considers the molecule structural similarity between and , while considers the scaffold structural similarity between and . Another issue with the conventional triplet loss is that it imposes no constraint on the representation space beyond the margin. It means that the actual representation distances are not necessarily to be well correlated with the computed margin, even if the margin constraints are fully satisfied. This is further elaborated by an example in Supplementary Information. Therefore, we include a secondary term into the adaptive margin loss by considering the structural relationship among the anchor and different negative samples. Detailed formulation of the adaptive margin loss is included in Method section. With careful consideration of existing structural similarity, the learned representation space would better capture molecule relationships in a fine-grained representation space.
@@
2.2 Molecular property prediction
To demonstrate the effectiveness of our approach, we first evaluate it on seven challenging classification datasets from MoleculeNet [56], which is a large-scale curated benchmark that covers multiple molecular property domains (e.g., physiology, biophysics). The scaffold split scheme [23] is applied. The performance is evaluated using the ROC-AUC value. Each experimental result is averaged over three different runs following the prior works [52, 38, 13]. To demonstrate the effectiveness of our framework across different model architectures, we pre-train both a graph neural network GIN [59] and a graph transformer GPS [36], termed Ours and Ours, respectively. We compare our methods with eleven competitive molecular representation learning baselines, which cover a wide variety of pre-train SSL techniques, model architectures, and input representations (i.e., sequence, graph, 3d geometry). Specifically, we first consider five contrastive baselines, including GraphLoG [60], D-SLA [25], MolCLR [52], GraphMVP [28], and KANO [14]. These methods also share the commonality of using GNNs to encode molecular graphs. Besides, we include six predictive baselines, including Hu et. al [22], GROVER [38], MoLFormer [39], KPGT [27], GEM [13], and Uni-Mol [65]. MoLFormer serves as a strong baseline for sequence-based models, while GEM and Uni-Mol are strong baselines for 3D geometry models. Last but not least, we also train the GIN and GPS models from scratch. Table 1 shows the performance comparison results. Our methods improve over the no-pretrain setting (i.e., GIN and GPS) by 12.6% and 8% ROC-AUC in average, respectively. When compared to the other SSL methods, our approach reaches the new state-of-the-art performance on BBBP, Clintox, BACE and SIDER datasets, while remaining highly competitive in the rest of the tasks. Our overall ROC-AUC score is 0.8% higher than the second best method (Uni-Mol). Notably, Uni-Mol leverages 3D geometric information and is an order of magnitude larger (in term of model parameters) than both Ours and Ours.
2.3 Binding potency prediction
@@
Figure 2a shows the average performance of thirteen methods on the MoleculeACE benchmark under stratified sampling [50]. Detailed table is included in Supplementary Information. The multi-layer perceptron (MLP) is trained using the ECFP4 fingerprint. We also highlight the differences in model sizes (i.e., number of parameters), as indicated by the x-axis and the size of the dots, and types of input representations (blue=1D sequence, red=2D topological graph, and yellow = 3D geometry). In general, Ours ranks in the second place, but with a significantly smaller model size than KPGT. Besides, almost all methods fail to surpass MLP with fingerprints. MolCLR and GraphLoG also show negative transfer compared to the GIN model. It demonstrates the incompetence of existing methods in learning the nuances of chemical knowledge behind the structural-property relationship, regardless of the input representations, pre-training strategies, and the model sizes [50, 9]. One of the reasons behind KPGT’s strong performance could be attributed to the fact that the molecule fingerprint and descriptors are heavily embedded within the model. In addition, KPGT is pre-trained using the ChEMBL database [16], such that over 99% of the testing molecules in MoleculeACE are already exposed to the model during the pre-training stage. In contrast, our models only have seen less than 5% of the testing molecules during the pre-training stage. The plot also suggests that there are no clear advantages to using either 1D sequence or 3D geometry for binding potency prediction, especially when comparing the performance between Uni-Mol and MoLFormer.
We further study the model performance in relation to the presence of activity cliffs in each dataset, which is measured by the roughness index (ROGI) [2] with respect to the ECFP4 fingerprints and potency labels. A smaller ROGI value indicates a smoother structure-property landscape, hence less activity cliffs. Since ROGI is a quantitative SPR (QSPR) metric that analyzes the overall chemical space, it does not account for any distribution shift between chemical spaces. Hence, its correlation with model performance would be easily obscured by any distribution shift among training, validation, and testing sets. Therefore, this analysis is performed under random splits, following the work in [2]. As shown in Figure 2b, the model performance is negatively correlated with roughness of the landscape as expected [2]. Compared to the MLP with fingerprint and KPGT, our method is shown to be more robust against tasks with rough structure-property landscapes. In addition, the plot also indicates that our method performs better under data scarcity (the size of dots represents the size of datasets).
2.4 Representation robustness
@@
Our composed representation yields the lowest ROGI value to start with. In other words, our pre-trained knowledge is more transferable to the target application. This is also shown by our rapid convergence rate in the validation set, reaching a validation R-squared of 0.676 at epoch 10.
@@
Our representation preserves the chemical knowledge learned from pre-training better than others, leading to less overfitting and better robustness. Since representations are optimized towards the property labels, the Rand index drops continuously for all methods. It means that the encoded information gradually shifts from being structure-oriented to label-oriented. However, our method has the lowest drop in Rand index of 0.128, compared to the drop of 0.141 by GraphLoG and 0.157 by MolCLR. The visualization also shows that the representations of MolCLR begin to overfit to the labels starting from epoch 10, resulting in the loss of substantial structural relationships between molecules. This explains its low ROGI value and validation R-squared along the training process.
@@
Our representation seems to exhibit a better understanding of the nuances of chemical knowledge in activity cliffs. Our average cliff-noncliff distance ratio in the validation set is always above one and larger than that of GraphLoG and MolCLR. As illustrated by the triplet samples, the red arrow is longer than the blue arrow across fine-tuning epochs, while the closeness of cliff and non-cliff pairs in space (i.e., structurally similar) is maintained. For MolCLR, even though the red arrow can be much longer than the blue one, the molecules are distant away. It means that MolCLR fails to capture the structural similarity aspect of the activity cliffs.
@@
To further understand why our composed representation is more resilient to the representation shift and can better preserve pre-trained knowledge, we present more analysis on three diverse datasets, along with the individual channel-wise representation behavior during fine-tuning. We choose CHEMBL237_Ki and CHEMBL262_Ki as the representative regression-based datasets of different scales, and BBBP as the typical classification-based dataset. The Rand index of representation clustering difference between the initial and the current timestamp is computed as a proxy (slightly different than before) for the representation shift. A smaller Rand index indicates a larger shift. As illustrated in Figure 4, channels with the highest prompt weights often exhibit the largest shift, and vice versa. This is reasonable because these channels contribute more to the optimization against the labels. Conversely, even though low-weighted channels contribute less to the optimization, they are more likely to preserve the pre-trained knowledge. As a result, the composed representation derived from channel aggregation exhibits a certain level of resilience to representation shifts, making it potentially more robust than other methods. We also include a few-shot representation probing analysis in Supplementary Information, which demonstrates similar patterns.
2.5 Activity cliffs analysis
@@
As shown in Figure 5, two series of compounds sharing the same scaffolds are potential inhibitors of glycogen synthase kinase-3 beta (GSK3). The molecule activity cliff pairs, determined by the formulation in [50], are compounds <>, <>, <>, and <>. The explanations of both our and MolCLR’s predictions are compared. In Figure 5a, the potential intra-molecular halogen-bonding contact between the chlorine atom and hydrazone in compound , as indicated by the green dashed line, is disfavored for the inter-molecular hydrogen-bonding contact between the backbone carbonyl of the active site VAL-135 and hydrazone of the compound [3]. As shown by our model explanation, compared to the compound and , the chlorine atom of compound , along with its associated benzene ring, contribute less to the overall binding affinity prediction. It aligns well with the predicted binding mode.
@@
3 Discussion
In this work, we propose a novel multi-channel learning framework for molecular representation learning, which aims to encode and utilize robust chemical knowledge that is generalizable to diverse downstream applications. Each channel is dedicated to learning unique self-supervised tasks, focusing on distinct yet correlated global and local aspects of the molecule. Specifically, MCD captures the molecule’s global similarity, SCD highlights the fundamental role of the scaffold in affecting molecular characteristics, and CP targets the composition of functional groups. During fine-tuning, the model is able to identify which channel-wise representation is most relevant to the current application, thereby making the composite representation context-dependent.
One limitation of this framework is the need for a more effective prompt weight optimization mechanism. The initialization of prompt weights using roughness index can lead to sub-optimal performance. Since roughness index is a global QSPR metric that targets the overall chemical space, it does not account for any distribution shift between training and testing sets. This is the same for the other QSPR measures as well (e.g., SALI [20], SARI [32]). As a result, the final representation performance may be less correlated with the initial roughness value. This also explains the performance gap between Figure 2a and Figure 2b.
There are several interesting directions for future research. One promising direction is to incorporate different input representations into the framework. By merely leveraging topological molecular structure, the model is unable to differentiate molecular components with different conformations (e.g., functional groups’ orientation or atom’s chirality), which could significantly alter biochemical behaviors. Besides, there exist other advanced data-driven techniques for studying the structural-activity relationship (SAR) that might be compatible with our framework. For example, Molecular Anatomy [30] argues that the network clustering from scaffold fragmentation and abstraction allows high quality SAR analysis. Such investigations aim to transfer knowledge from cheminformatics to machine learning models, potentially improving both model interpretability and robustness. More importantly, while our method has immediate implications for drug discovery, its molecular representation robustness further shed lights on its promising potential in other sub-fields of chemistry, such as materials science and environmental chemistry.
4 Methods

### v3

# Multi-channel learning for integrating structural hierarchies into context-dependent molecular representation

@@ Abstract
Reliable molecular property prediction is essential for various scientific endeavors and industrial applications, such as drug discovery. However, the data scarcity, combined with the highly non-linear causal relationships between physicochemical and biological properties and conventional molecular featurization schemes, complicates the development of robust molecular machine learning models. Self-supervised learning (SSL) has emerged as a popular solution, utilizing large-scale, unannotated molecular data to learn a foundational representation of chemical space that might be advantageous for downstream tasks. Yet, existing molecular SSL methods largely overlook chemical knowledge, including molecular structure similarity, scaffold composition, and the context-dependent aspects of molecular properties when operating over the chemical space. They also struggle to learn the subtle variations in structure-activity relationship. This paper introduces a multi-channel pre-training framework that learns robust and generalizable chemical knowledge. It leverages the structural hierarchy within the molecule, embeds them through distinct pre-training tasks across channels, and aggregates channel information in a task-specific manner during fine-tuning. Our approach demonstrates competitive performance across various molecular property benchmarks and offers strong advantages in particularly challenging yet ubiquitous scenarios like activity cliffs.
Introduction
@@
To approach the aforementioned challenges, we introduce a prompt-guided multi-channel learning framework for molecular representation learning. Each of the channels, guided by a specific prompt token, is responsible for learning one dedicated SSL task. Essentially, the pre-trained model is able to learn distinct representation spaces. During fine-tuning, a prompt selection module aggregates representations into a composite representation and uses it for the downstream molecular property predictions. This involves determining which information channel is most relevant to the current application, thereby making the representation context-dependent. We later show how this composite formulation is more resilient to label overfitting and manifests better robustness. In addition, we design the pre-train tasks to form an interpolation from a global view to a local view of the molecular structures. Besides leveraging the global molecule contrastive learning and the local context prediction [19], we introduce the task of scaffold contrastive distancing, highlighting the fundamental role of scaffolds in affecting molecular characteristics and behaviors. Since scaffolds are often treated as starting points for new compound design, scaffold distancing aims to map molecules with similar scaffolds (generated via scaffold-invariant perturbations) closer in the representation space. Additionally, it pushes molecules with different scaffolds apart, where the distance margin is computed adaptively based on structure composition difference. Note that scaffold distancing tackles the partial but core view of the molecules. The overall framework is pre-trained using ZINC15 [38], and evaluated on 7 molecular property prediction tasks in MoleculeNet [5] and 30 binding potency prediction tasks in MoleculeACE [6]. Learning to leverage information from different channels for different applications, our method surpasses various representation learning baselines in both benchmarks. More importantly, our method is shown to handle the challenge of activity cliffs more effectively, whereas competing approaches are more susceptible to the negative transfer, leading to a substantial performance decline. This suggests that these methods may rely more on surface-level patterns even after pre-training or are more susceptible to knowledge forgetting during fine-tuning, causing them to struggle with challenging problems that require a nuanced understanding of chemical knowledge. On the contrary, our learned representation demonstrates enhanced ability in preserving pre-trained knowledge during fine-tuning, offering improved transferability and robustness compared to other baselines. Case study shows that our method has the potential to identify crucial patterns that contribute to activity cliffs, even when relying solely on topological information.
Results
@@
The proposed framework (Figure 1) comprises three major components that differ from the conventional pretrain-finetune paradigm on molecules: (1) The prompt-guided multi-channel learning, (2) contrastive learning with adaptive margin, and (3) scaffold-invariant molecule perturbation. It demonstrates effectiveness on both the molecular property prediction [5] and binding potency prediction [6] benchmarks, offering enhanced robustness and interpretability.
@@
We now briefly introduce the three learning channels. (i) Molecule distancing (Figure 1b) is archived using a variant of triplet contrastive loss [40]. It considers triplet of molecule samples {anchor, positive, negative}, where negative (i.e., dissimilar) samples are pushed apart against the anchor and positive (i.e., similar) samples by a margin . On top of this, we propose the adaptive margin , as detailed below. It introduces another level of distancing constraints based on the structural similarity of molecule composition. We follow the work of Molecular Contrastive Learning of Representations (MolCLR) [21] and generate positive samples via molecule subgraph masking. (ii) Scaffold distancing (Figure 1c) is a contrastive learning task that focuses on scaffold differences. Molecule scaffolds are viewed as the foundation for a range of biologically active molecules. They play a crucial role in drug discovery and medicinal chemistry by providing a starting point for compound designs with desired pharmacological properties. In other words, molecules with similar scaffolds are more likely to possess similar physical (e.g., solubility, lipophilicity) and biological (e.g., conformational property when interacting with a protein) characteristics, and thereby sharing similar semantics. Scaffold distancing contrasts the scaffold-invariant molecule perturbations, as detailed below, against molecules with different scaffolds using the adaptive margin loss. (iii) Context prediction (Figure 1d) involves masked subgraph prediction and motif prediction, which are also adopted in GROVER [19]. For each molecular graph, a random subgraph (i.e., a center atom and its one-hop neighbors) is masked out, and the model aims to reconstruct the subgraph based on its surrounding structures. Motif prediction aims to predict the existence of functional groups within the molecule. This learning channel mainly focuses on the local view of the molecule by identifying the existence of substructure and functional groups, while molecule distancing and scaffold distancing focus more on the global view and partial view, respectively. Each channel has its own readout layer, termed prompt-guided node aggregation (Figure 1e). Presumably, the distribution of atom importance should be different across channels (Supplementary Figs. 7-9 and Supplementary Note 6). The learned channel-wise representations exhibit a high correspondence to the associated structural features (Supplementary Fig. 1 and Supplementary Note 1) and support hierarchical iterative clustering of the chemical space (Supplementary Fig. 2 and Supplementary Note 2). To further improve the robustness of the learned knowledge under this framework, we incorporate two regularization tricks on the intra-channel node aggregation and the inter-channel alignment. For the molecule and scaffold channels, the aggregation attentions are more encouraged to span across all atoms and scaffold atoms, respectively. Meanwhile, a small set of supervised tasks is utilized to regularize the composite representation under three prompt weight presets, hence improving the channel alignment. A comprehensive ablation study is provided in Supplementary Fig. 10 and Supplementary Note 7.1.
@@
In the fine-tuning stage, the model is initialized with the same molecule encoder and prompt-guided aggregation modules, along with the learned parameters from pre-training. The parameters in the aggregation modules are fixed during fine-tuning as we aim to use it as a pooling layer independent of the downstream applications. In addition, a prompt-tuning module is introduced to determine which channel is most relevant to the current application. It essentially learns a task-specific prompt distribution. The prompt weights are utilized to linearly combine the channel-wise information into a composite molecule representation, which is then used for the task-specific prediction. We discover that this approach is more effective than simply concatenating the representations (Supplementary Fig. 11 and Supplementary Note 7.2). We initialize the prompt weights by choosing the candidate which leads to the smoothest quantitative structure-property landscape (i.e., lowest roughness index) [41] of the composite representation. More details are provided in the Method section.
@@
We further introduce the adaptive margin loss, a variant of the triplet loss [40], that supports the contrastive learning in the first two channels (molecule distancing and scaffold distancing). In the conventional triplet loss, the representation distance between the anchor and negative (i.e., semantic-dissimilar) sample needs to be at least by margin larger than the distance between the anchor and positive (i.e., semantic similar) sample . Note that this margin remains the same for any triplet considered. However, when applied to molecule triplets, it neglects the known structural relationship between molecules (e.g., co-existence of functional groups). To learn a more fine-grained molecule representation space, we propose to adaptively compute the molecule triplet margin based on the Tanimoto similarity between molecule fingerprints. As shown in Figure 1b and c, the adaptive margin considers the molecule structural similarity between and , while considers the scaffold structural similarity between and . Another issue with the conventional triplet loss is that it imposes no constraint on the representation space beyond the margin. It means that the actual representation distances are not necessarily to be well correlated with the computed margin, even if the margin constraints are fully satisfied. This is further elaborated the example in Supplementary Fig. 3. Therefore, we include a secondary term into the adaptive margin loss by considering the structural relationship among the anchor and different negative samples. Detailed formulation of the adaptive margin loss is included in Method section. With careful consideration of existing structural similarity, the learned representation space would better capture molecule relationships in a fine-grained representation space. A performance drop is observed when the adaptive margin loss is replaced with the conventional margin loss (Supplementary Fig. 10 and Supplementary Note 7.1).
@@
Molecular property prediction
To demonstrate the effectiveness of our approach, we first evaluate it on seven challenging classification datasets from MoleculeNet [5], which is a large-scale curated benchmark that covers multiple molecular property domains (e.g., physiology, biophysics). The scaffold split scheme [42] is applied. The performance is evaluated using the ROC-AUC value. Each experimental result is averaged over three different runs following the prior works [21, 19, 25]. To demonstrate the effectiveness of our framework across different model architectures, we pre-train both a graph neural network GIN [43] and a graph transformer GPS [44], termed Ours and Ours, respectively. We compare our methods with twelve competitive molecular representation learning baselines, which cover a wide variety of pre-train SSL techniques, model architectures, and input representations (i.e., sequence, graph, geometry). Specifically, we first consider four contrastive baselines, including GraphLoG [22], D-SLA [23], MolCLR [21], and KANO [24]. These methods also share the commonality of using GNNs to encode molecular graphs. Besides, we include six predictive baselines, including Hu et. al [18], GROVER [19], MoLFormer [28], KPGT [45], GEM [25], and Uni-Mol [27]. MoLFormer serves as a strong baseline for sequence-based models, while GEM and Uni-Mol are strong baselines for 3D geometry models. We further incorporate GraphMVP [46] and ImageMol [47] as two multi-task learning baselines whose pre-training tasks cover both contrastive and predictive learning. Last but not least, we train the GIN and GPS models from scratch. Table 1 shows the performance comparison results. Our methods improve over the no-pretrain setting (i.e., GIN and GPS) by 12.6% and 8% ROC-AUC in average, respectively. When compared to the other SSL methods, our approach reaches the new state-of-the-art performance on BBBP, Clintox, BACE and SIDER datasets, while remaining highly competitive in the rest of the tasks. Our overall ROC-AUC score is 0.8% higher than the second best method (Uni-Mol). Notably, Uni-Mol leverages 3D geometric information and is an order of magnitude larger (in term of model parameters) than both Ours and Ours.
Binding potency prediction
@@
Figure 2a shows the performance of thirteen methods on the MoleculeACE benchmark under stratified splits [6] (Supplementary Table 2). The multi-layer perceptron (MLP) is trained using the ECFP4 fingerprint. We also compare the average performance with respect to the model sizes (i.e., number of parameters) in Figure 2b, as indicated by its x-axis and the size of the dots, and types of input representations (blue = 1D sequence, red = 2D topological graph, and yellow = 3D geometry). In general, Ours ranks in the second place, but with a significantly smaller model size than KPGT. Besides, almost all methods fail to surpass MLP with fingerprints. MolCLR and GraphLoG also show negative transfer compared to the GIN model. It demonstrates the incompetence of existing methods in learning the nuances of chemical knowledge behind the structural-property relationship, regardless of the input representations, pre-training strategies, and the model sizes [6, 12]. One of the reasons behind KPGT’s strong performance could be attributed to the fact that the molecule fingerprint and descriptors are heavily embedded within the model. In addition, KPGT is pre-trained using the ChEMBL database [48], such that over 99% of the testing molecules in MoleculeACE are already exposed to the model during the pre-training stage. In contrast, our models only have seen less than 5% of the testing molecules during the pre-training stage. The plot also suggests that there are no clear advantages of molecular representation learning using either 1D sequence or 3D geometry, especially when comparing the performance with Uni-Mol and MoLFormer.
We further study the model performance in relation to the presence of activity cliffs in each dataset, which is measured by the roughness index (ROGI) [41] with respect to the ECFP4 fingerprints and potency labels. A smaller ROGI value indicates a smoother structure-property landscape, hence less activity cliffs. Since ROGI is a quantitative SPR (QSPR) metric that analyzes the overall chemical space, it does not account for any distribution shift between chemical spaces. Hence, its correlation with model performance would be easily obscured by any distribution shift among training, validation, and testing sets (Supplementary Fig. 14). For this reason, all experiments in this work that involves QSPR analysis are performed using random splits, following the work in [41]. As shown in Figure 2c and Supplementary Table 3, the model performance is negatively correlated with landscape roughness for all methods, which is consistent with the results in [41]. Compared to the MLP with fingerprint and KPGT, our method is shown to be more robust against tasks with rough structure-property landscapes. In addition, the plot also indicates that our method performs better under data scarcity, where the size of dots represents the size of datasets.
Representation robustness
@@
Our composed representation yields the lowest ROGI value to start with. In other words, our pre-trained knowledge is more transferable to the target application. This is also shown by our rapid convergence rate in the validation set, reaching a validation R-squared of 0.676 at epoch 10 (Supplementary Fig. 5 and Supplementary Note 4).
@@
Our representation preserves the chemical knowledge learned from pre-training better than others, leading to less overfitting and better robustness. Since representations are optimized towards the property labels, the Rand index drops continuously for all methods. It means that the encoded information gradually shifts from being structure-oriented to label-oriented. However, our method has the lowest drop in Rand index of 0.072, compared to the drop of 0.09 by GraphLoG and 0.181 by MolCLR. The visualization also shows that the representations of MolCLR begin to overfit to the labels starting from epoch 10, resulting in the loss of substantial structural relationships between molecules. This explains its low ROGI value and validation R-squared along the training process.
@@
Our representation seems to exhibit a better understanding of the nuances of chemical knowledge in activity cliffs. Our average cliff-noncliff distance ratio in the validation set is always above one and larger than that of GraphLoG and MolCLR. As illustrated by the triplet samples, the red arrow is longer than the blue arrow across fine-tuning epochs, while the closeness of cliff and non-cliff pairs in space (i.e., structurally similar) is maintained. For MolCLR, even though the red arrow can be much longer than the blue one, the molecules are distant away. It means that MolCLR fails to capture the structural similarity aspect of the activity cliffs. A detailed examination of the distance histogram is visualized in Supplementary Fig. 13.
@@
To further understand why our composed representation is more resilient to the representation shift and can better preserve pre-trained knowledge, we present more analysis on three diverse datasets, along with the individual channel-wise representation behavior during fine-tuning. We choose CHEMBL237_Ki and CHEMBL262_Ki as the representative regression-based datasets of different scales, and BBBP as the typical classification-based dataset. The Rand index of representation clustering difference between the initial and the current timestamp is computed as a proxy (slightly different than before) for the representation shift. A smaller Rand index indicates a larger shift. As illustrated in Figure 4, channels with the highest prompt weights often exhibit the largest shift, and vice versa. This is reasonable because these channels contribute more to the optimization against the labels. Conversely, even though low-weighted channels contribute less to the optimization, they are more likely to preserve the pre-trained knowledge. As a result, the composed representation derived from channel aggregation exhibits a certain level of resilience to representation shifts, making it potentially more robust than other methods. The probing analysis under the few-shot learning settings also reveals similar patterns (Supplementary Fig. 6 and Supplementary Note 5).
Activity cliffs analysis
@@
As shown in Figure 5, visualized by PyMOL [52], two series of compounds sharing the same scaffolds are potential inhibitors of glycogen synthase kinase-3 beta (GSK3). The molecule activity cliff pairs, determined by the formulation in [6], are compounds <>, <>, <>, and <>. The explanations of both our and MolCLR’s predictions are compared. In Figure 5a, the potential intra-molecular halogen-bonding contact between the chlorine atom and hydrazone in compound , as indicated by the green dashed line, is disfavored for the inter-molecular hydrogen-bonding contact between the backbone carbonyl of the active site VAL-135 and hydrazone of the compound [53]. As shown by our model explanation, compared to the compound and , the chlorine atom of compound , along with its associated benzene ring, contribute less to the overall binding affinity prediction. It aligns well with the predicted binding mode.
@@
Discussion
In this work, we propose a multi-channel learning framework for molecular representation learning, which aims to encode and utilize robust chemical knowledge that is generalizable to diverse downstream applications. Each channel is dedicated to learning unique self-supervised tasks, focusing on distinct yet correlated global and local aspects of the molecule. Specifically, MCD captures the molecule’s global similarity, SCD highlights the fundamental role of the scaffold in affecting molecular characteristics, and CP targets the composition of functional groups. During fine-tuning, the model is able to identify which channel-wise representation is most relevant to the current application, thereby making the composite representation context-dependent.
One limitation of this framework is the need for a more effective prompt weight optimization mechanism. The initialization of prompt weights using roughness index can lead to sub-optimal performance. Since roughness index is a global QSPR metric that targets the overall chemical space, it does not account for any distribution shift between training and testing sets. This is the same for the other QSPR measures as well (e.g., SALI [56], SARI [57]). As a result, the final representation performance may be less correlated with the initial roughness value under designated splits. This also explains the performance gap between Figure 2a and Figure 2b.
There are several interesting directions for future research. One promising direction is to incorporate different input representations into the framework. By merely leveraging topological molecular structure, the model is unable to differentiate molecular components with different conformations (e.g., functional groups’ orientation or atom’s chirality), which could significantly alter biochemical behaviors (see Supplementary Note 9 for further discussion). Besides, there exist other advanced data-driven techniques for studying the structural-activity relationship (SAR) that might be compatible with our framework. For example, Molecular Anatomy [58] argues that the network clustering from scaffold fragmentation and abstraction allows high quality SAR analysis. Such investigations aim to transfer knowledge from cheminformatics to machine learning models, potentially improving both model interpretability and robustness. More importantly, while our method has immediate implications for drug discovery, its molecular representation robustness further shed lights on its promising potential in other sub-fields of chemistry, such as materials science and environmental chemistry.
Methods